OpenStudy (anonymous):

Help Please!!! Frank mixed two different cereals to make 10 lb. cereal bags. If brand A costs $2.10/lb. and brand B costs $2.60/lb. how many pounds of each brand should he mix to sell the bags $25.

jimthompson5910 (jim_thompson5910):

Let x = amount of brand A cereal (in pounds) and y = amount of brand B cereal (in pounds) The word problem translates to the two equations x+y = 10 2.10x + 2.60y = 25*10

jimthompson5910 (jim_thompson5910):

use either substitution or elimination to solve for x and y.